# The Moroccan Public Procurement Game

### v1

@@ 5 The $N$-players version
$g_{i}(x_{1},\ldots,x_{N})$ $=\sum_{n=0}^{N-1}\sum_{J_{n}(i)}\frac{1}{n+1}\prod_{k\in J_{n}(i)}\mathbf{1}_{\{x_{i}=x_{k}\}}\left(\mathbf{1}_{\{x_{i}\leqslant P\}}\prod_{j\neq i,\,j\in N^{-}}\mathbf{1}_{\{x_{j}<x_{i}\}}+\mathbf{1}_{\{x_{i}>P\}}\prod_{j\not\in J_{n}(i),\,j=1}^{N}\mathbf{1}_{\{x_{i}<x_{j}\}}\right)$ (77)
@@ 5.2 Existence result for $N\geqslant 3$
Thus, the game is better-reply secure (Reny1999). However, the game is not quasiconcave which prevent the application of Reny’s fundamental result about equilibrium existence in pure strategies.
@@
On can check that our game is not uniformly payoff secure, it is not uniformly diagonally secure in the sens of Prokopovych2014 neither.
@@
We now construct such a sequence of deviations $(x_{i}^{k})$ for our game $\{(g_{i},[A,B])_{i=1}^{N}\}$. For any $x_{i}\in[A,B]$, the set of discontinuities $\mathbb{D}_{i}(x_{i})$ is the union of the following hypersurfaces:
@@
However, we get a problem at ties situation: If $\mathbf{x}_{-i}\in\mathbb{T}_{i}(x_{i})$, one can check $x_{i}^{k}\uparrow x_{i}$ is not a good choice when $x_{i}<E$. To bypass this obstacle, we define the game $\{(\tilde{g}_{i},[A,B])_{i=1}^{N}\}$ where each payoff function $\tilde{g}_{i}$ has zero value on the ties. One can check for $x_{i}^{k}\downarrow x_{i}$ and $\mathbf{x}_{-i}\in\mathbb{T}_{i}(x_{i})$ that
@@
To transfer the diagonal payoff security to the original game $\{(g_{i},[A,B])_{i=1}^{N}\}$ we will first need some a measure smooth approximation, we report the reader to the appendix for the proof.
Lemma 5.4.
Thin-mass approximation on a diagonal hypersurface
Let $S_{i}\subset{\mathbb{R}}$ be compact for $i=1,\dots,N$, and write $S=\prod_{i=1}^{N}S_{i}$ with the product Borel sigma-algebra. Let $\mathcal{D}\subset S$ be a Borel set which is a finite union of embedded hypersurfaces in $\mathbb{R}$} (in particular $\mathcal{D}$ has Lebesgue measure zero). Then for every product probability measure $\boldsymbol{\sigma}=(\sigma_{1},\ldots,\sigma_{N})\in\prod_{i=1}^{N}\Delta(S_{i})$ there exists a product measure $\boldsymbol{\sigma}^{\prime}=(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{N})\in\prod_{i=1}^{N}\Delta(S_{i})$ such that
$\text{(i)}\quad\boldsymbol{\sigma}^{\prime}\xrightarrow{w}\boldsymbol{\sigma}\quad\text{(weak convergence)},\qquad\text{(ii)}\quad\boldsymbol{\sigma}^{\prime}(\mathcal{D})=0.$ (100)
In particular, given any weak neighbourhood $\mathcal{U}$ of $\boldsymbol{\sigma}$ one can choose $\boldsymbol{\sigma}^{\prime}\in\mathcal{U}$ satisfying $\boldsymbol{\sigma}^{\prime}(\mathcal{D})=0$.
We will now prove that the original game $\{(g_{i},[A,B])_{i=1}^{N}\}$ is diagonally payoff secure, that is, $\forall\mu\in\Delta([A,B])$, $\boldsymbol{\mu}=\underset{N\,\text{times}}{\underbrace{(\mu,\ldots,\mu)}}$, there exists $\tilde{x}\in[A,B]$ such that,
$\int g_{i}\,d\boldsymbol{\mu}-\int g_{i}(x,\ldots,\tilde{x},\ldots,x)\,d\boldsymbol{\lambda}_{-i}$ $\leq\varepsilon.$ (101)
for all $\lambda\in N(\mu)$, $\boldsymbol{\lambda}_{-i}=\underset{(N-1)\,\text{times}}{\underbrace{(\lambda,\ldots,\lambda)}}$.
Define the following measure-smooth approximations as in Lemma 5.4:
$\text{(i)}\quad\boldsymbol{\mu}^{\prime}\xrightarrow{w}\boldsymbol{\mu},\qquad\text{(ii)}\quad\boldsymbol{\mu}^{\prime}(\mathcal{D}_{i})=0.$
$\text{(i)}\quad\boldsymbol{\lambda}^{\prime}\xrightarrow{w}\boldsymbol{\lambda},\qquad\text{(ii)}\quad\boldsymbol{\lambda}^{\prime}(\mathcal{D}_{i})=0.$
In particular $\boldsymbol{\mu}^{\prime}(\mathcal{T}_{i})=0$ and $\boldsymbol{\lambda}^{\prime}(\mathcal{T}_{i})=0$. By diagonal payoff security of the game $\{(\tilde{g}_{i},[A,B])_{i=1}^{N}\}$, for $\boldsymbol{\mu}^{\prime}$, there exists $\bar{x}\in[A,B]$ (the point that we gonna choose for to prove the inequality) such that,
$\int\tilde{g}_{i}\,d\boldsymbol{\mu}^{\prime}-\int\tilde{g}_{i}(x,\ldots,\bar{x},\ldots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}$ $\leq\frac{\varepsilon}{5}.$ (102)
for all $\lambda^{\prime}\in N(\mu^{\prime})$.
Choosing $\tilde{x}=\bar{x}$, we apply the triangle inequality to obtain
$\begin{split}\int g_{i}\,d\boldsymbol{\mu}\;-\;\int g_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}_{-i}&\leq\Big|\,\int g_{i}\,d\boldsymbol{\mu}-\int g_{i}\,d\boldsymbol{\mu}^{\prime}\,\Big|+\Big|\,\int g_{i}\,d\boldsymbol{\mu^{\prime}}-\int\tilde{g}_{i}\,d\boldsymbol{\mu^{\prime}}\,\Big|\\ &+\Big|\,\int\tilde{g}_{i}\,d\boldsymbol{\mu^{\prime}}-\int\tilde{g}_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}\,\Big|\\ &+\Big|\,\int\tilde{g}_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}-\int g_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}\,\Big|\\ &+\Big|\,\int g_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}-\int g_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}_{-i}\,\Big|.\end{split}$ (103)
We now bound each term.
5.2.0.1 (1) Approximation of $\boldsymbol{\mu}$.
Fix the (arbitrary) product measure $\boldsymbol{\mu}$ under consideration and let $\delta:=\varepsilon/9$. Since $g_{i}$ is bounded on the compact space $A,B]^{N}$, Lusin’s theorem implies that there exists a closed set $K\subset[A,B]^{N}$ such that $g_{i}$ is continuous on $K$ and
$\boldsymbol{\mu}(K)>1-\delta.$
Set $U:=[A,B]^{N}\setminus K$; then $U$ is open and $\boldsymbol{\mu}(U)<\delta$. By Lemma 5.4 we may choose the approximating product measure $\boldsymbol{\mu}^{\prime}$ arbitrarily close to $\boldsymbol{\mu}$ in the weak topology and in particular so that $\boldsymbol{\mu}^{\prime}(U)<\delta$ and $\boldsymbol{\mu}^{\prime}\xrightarrow{w}\boldsymbol{\mu}$.
Decompose
$\Big|\int g_{i}\,d\boldsymbol{\mu}-\int g_{i}\,d\boldsymbol{\mu}^{\prime}\Big|\leq\Big|\int_{K}g_{i}\,d\boldsymbol{\mu}-\int_{K}g_{i}\,d\boldsymbol{\mu}^{\prime}\Big|+\boldsymbol{\mu}(U)+\boldsymbol{\mu}^{\prime}(U).$ (104)
On the compact set $K$ the function $g_{i}$ is continuous and bounded, hence weak convergence $\boldsymbol{\mu}^{\prime}\to\boldsymbol{\mu}$ implies that the first term can be made $\delta$ for $\boldsymbol{\mu}^{\prime}$ sufficiently close to $\boldsymbol{\mu}$. Using $\boldsymbol{\mu}(U)<\delta$ and $\boldsymbol{\mu}^{\prime}(U)<\delta$ we obtain
$\Big|\int g_{i}\,d\boldsymbol{\mu}-\int g_{i}\,d\boldsymbol{\mu}^{\prime}\Big|<3\delta=\frac{\varepsilon}{3}.$ (105)
5.2.0.2 (2) Elimination of the tie set under $\boldsymbol{\mu}^{\prime}$.
Because $\boldsymbol{\mu}^{\prime}(\mathcal{T}_{i})=0$ and $g_{i}=\tilde{g}_{i}$ on $A,B]^{N}\setminus\mathcal{T}$,
$\Big|\,\int g_{i}\,d\boldsymbol{\mu^{\prime}}-\int\tilde{g}_{i}\,d\boldsymbol{\mu^{\prime}}\,\Big|=0.$ (106)
5.2.0.3 (3) Diagonal payoff security for the tie-broken game.
From diagonal payoff security of $\tilde{g}_{i}$ at $\boldsymbol{\mu}^{\prime}$ we have
$\Big|\,\int\tilde{g}_{i}\,d\boldsymbol{\mu^{\prime}}-\int\tilde{g}_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}\,\Big|\leq\frac{\varepsilon}{3}\quad\text{for all }\lambda^{\prime}\in N(\mu^{\prime}).$ (107)
5.2.0.4 (4) Elimination of the tie set under $\boldsymbol{\lambda}^{\prime}$.
Since $\boldsymbol{\lambda}^{\prime}(\mathcal{T}_{i})=0$ and $g_{i}=\tilde{g}_{i}$ outside $\mathcal{T}_{i}$,
$\Big|\,\int\tilde{g}_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}-\int g_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}^{\prime}_{-i}\,\Big|=0.$ (108)
5.2.0.5 (5) Approximation of $\boldsymbol{\lambda}_{-i}$.
Fix the (arbitrary) measure $\boldsymbol{\lambda}_{-i}=(\lambda,\ldots,\lambda)$ under consideration and set $\delta:=\varepsilon/9$. Consider the function
$h(x):=g_{i}(x,\dots,\tilde{x},\dots,x),$
viewed as a bounded measurable function on $A,B]^{\,N-1}$ with $0\leq h\leq 1$. By Lusin’s theorem there exists a closed set $K\subset[A,B]^{N-1}$ such that $h$ is continuous on $K$ and
$\boldsymbol{\lambda}_{-i}(K)>1-\delta.$
Set $U:=[A,B]^{N-1}\setminus K$; then $U$ is open and $\boldsymbol{\lambda}_{-i}(U)<\delta$.
By Lemma 5.4 we may choose the approximating product measure $\boldsymbol{\lambda}^{\prime}_{-i}$ arbitrarily close to $\boldsymbol{\lambda}_{-i}$ in the weak topology, and in particular so that $\boldsymbol{\lambda}^{\prime}_{-i}(U)<\delta$ and $\boldsymbol{\lambda}^{\prime}_{-i}\xrightarrow{w}\boldsymbol{\lambda}_{-i}$. Decompose
$\Big|\int h\,d\boldsymbol{\lambda}^{\prime}_{-i}-\int h\,d\boldsymbol{\lambda}_{-i}\Big|\leq\Big|\int_{K}h\,d\boldsymbol{\lambda}^{\prime}_{-i}-\int_{K}h\,d\boldsymbol{\lambda}_{-i}\Big|+\boldsymbol{\lambda}^{\prime}_{-i}(U)+\boldsymbol{\lambda}_{-i}(U).$ (109)
Since $h$ is continuous and bounded on the compact set $K$, weak convergence $\boldsymbol{\lambda}^{\prime}_{-i}\to\boldsymbol{\lambda}_{-i}$ implies the first term can be made $\delta$ for $\boldsymbol{\lambda}^{\prime}_{-i}$ sufficiently close to $\boldsymbol{\lambda}_{-i}$. Using $\boldsymbol{\lambda}_{-i}(U)<\delta$ and $\boldsymbol{\lambda}^{\prime}_{-i}(U)<\delta$ we obtain
$\Big|\int h\,d\boldsymbol{\lambda}^{\prime}_{-i}-\int h\,d\boldsymbol{\lambda}_{-i}\Big|<3\delta=\frac{\varepsilon}{3}.$ (110)
Thus by choosing $\boldsymbol{\lambda}^{\prime}_{-i}$ suitably close to $\boldsymbol{\lambda}_{-i}$ we achieve the desired bound.
Adding the five bounds 105, 106, 107, 108, and 110 gives
$\int g_{i}\,d\boldsymbol{\mu}-\int g_{i}(x,\dots,\tilde{x},\dots,x)\,d\boldsymbol{\lambda}_{-i}\;\leq\;\varepsilon.$ (111)
We just proved the following theorem.
@@
The game $\{(g_{i},[A,B])_{i=1}^{N}\}$ is diagonally payoff secure.
Diagonal payoff security together with reciprocal upper semicontinuity implies diagonal better–reply security Reny1999. One can therefore deduce the existence of a symmetric Nash equilibrium for our game.
Theorem 5.6.
If the game $\{(g_{i},S_{i})_{i=1}^{N}\}$ is quasi-symmetric, compact, diagonally quasiconcave, and diagonally better-reply secure, then it possesses a symmetric pure strategy Nash equilibrium.
@@
Lemma 5.9.
Thin-mass approximation on a diagonal hypersurface
Let $S_{i}\subset{\mathbb{R}}$ be compact for $i=1,\dots,N$, and write $S=\prod_{i=1}^{N}S_{i}$ with the product Borel sigma-algebra. Let $\mathcal{D}\subset S$ be a Borel set which is a finite union of embedded hypersurfaces in $\mathbb{R}$} (in particular $\mathcal{D}$ has Lebesgue measure zero). Then for every product probability measure $\boldsymbol{\sigma}=(\sigma_{1},\ldots,\sigma_{N})\in\prod_{i=1}^{N}\Delta(S_{i})$ there exists a product measure $\boldsymbol{\sigma}^{\prime}=(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{N})\in\prod_{i=1}^{N}\Delta(S_{i})$ such that
$\text{(i)}\quad\boldsymbol{\sigma}^{\prime}\xrightarrow{w}\boldsymbol{\sigma}\quad\text{(weak convergence)},\qquad\text{(ii)}\quad\boldsymbol{\sigma}^{\prime}(\mathcal{D})=0.$
In particular, given any weak neighbourhood $\mathcal{U}$ of $\boldsymbol{\sigma}$ one can choose $\boldsymbol{\sigma}^{\prime}\in\mathcal{U}$ satisfying $\boldsymbol{\sigma}^{\prime}(\mathcal{D})=0$.
Proof.
Since each $S_{i}\subset{\mathbb{R}}$ is compact, we may regard each marginal $\sigma_{i}$ as a probability measure on $\mathbb{R}$} supported on $S_{i}$ (extend by 0 outside $S_{i}$). Fix a nonnegative $C_{c}^{\infty}$ function $\phi:{\mathbb{R}}\to[0,\infty)$ supported in the unit ball with $\int_{{\mathbb{R}}}\phi(x)\,dx=1$ (a standard mollifier). For $\varepsilon>0$ define the scaled mollifier
$\phi_{\varepsilon}(x):=\varepsilon^{-1}\phi(x/\varepsilon).$
For each marginal $\sigma_{i}$ define the mollified measure $\sigma_{i}^{\varepsilon}$ by convolution with $\phi_{\varepsilon}$:
$\sigma_{i}^{\varepsilon}:=\sigma_{i}*\phi_{\varepsilon},$
i.e. for every Borel set $A\subset{\mathbb{R}}$,
$\sigma_{i}^{\varepsilon}(A)=\int_{\mathbb{R}}\bigg(\int_{A}\phi_{\varepsilon}(x-y)\,dx\bigg)\,d\sigma_{i}(y).$
Equivalently, $\sigma_{i}^{\varepsilon}$ has a continuous density
$p_{i}^{\varepsilon}(x)=\int_{\mathbb{R}^{d}}\phi_{\varepsilon}(x-y)\,d\sigma_{i}(y),$
with respect to Lebesgue measure; in particular $\sigma_{i}^{\varepsilon}$ is absolutely continuous and supported in a small neighbourhood of $S_{i}$. Standard properties of mollification yield:
  1. 1.
    (Weak convergence) $\sigma_{i}^{\varepsilon}\xrightarrow{w}\sigma_{i}$ as $\varepsilon\downarrow 0$ for each $i$.
  2. 2.
    (Absolute continuity) for every $\varepsilon>0$ the measure $\sigma_{i}^{\varepsilon}$ is absolutely continuous w.r.t. Lebesgue measure and has a continuous density $p_{i}^{\varepsilon}$.
Define the measure $\boldsymbol{\sigma}^{\varepsilon}:=(\sigma_{1}^{\varepsilon},\ldots,\sigma_{N}^{\varepsilon})$. Then $\boldsymbol{\sigma}^{\varepsilon}$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}^{N}$ and has the continuous joint density
$p^{\varepsilon}(x_{1},\dots,x_{N})=\prod_{i=1}^{N}p_{i}^{\varepsilon}(x_{i}).$
Since $\mathcal{D}$ is a finite union of embedded hypersurfaces, it has Lebesgue measure zero in ${\mathbb{R}}^{N}$. Hence for every $\varepsilon>0$,
$\boldsymbol{\sigma}^{\varepsilon}(\mathcal{D})=0.$
Next we verify weak convergence of the product measures. By the weak convergence of marginals and standard properties of the product topology, we have
$\boldsymbol{\sigma}^{\varepsilon}\xrightarrow{w}\boldsymbol{\sigma}\qquad\text{as }\varepsilon\downarrow 0.$
which proves the lemma. ∎

### v2

@@ 5 The $N$-players version
$g_{i}(x_{1},\ldots,x_{N})$ $=\sum_{n=0}^{N-1}\sum_{J_{n}(i)}\frac{1}{n+1}\prod_{k\in J_{n}(i)}\mathbf{1}_{\{x_{i}=x_{k}\}}\left(\mathbf{1}_{\{x_{i}\leqslant P\}}\prod_{j\neq i,\,j\in N^{-}\setminus J_{n}(i)}\mathbf{1}_{\{x_{j}<x_{i}\}}+\mathbf{1}_{\{x_{i}>P\}}\prod_{j\neq i,\,j\not\in J_{n}(i)}\mathbf{1}_{\{x_{i}<x_{j}\}}\right)$ (77)
@@ 5.2 Existence result for $N\geqslant 3$
Thus, the game is better-reply secure (Reny1999, Prop. 3.2). However, the game is not quasiconcave which prevent the application of Reny’s fundamental result about equilibrium existence in pure strategies.
@@
On can check that our game is not uniformly payoff secure. Take for example $N=2$, and choose $x_{i}<E$, then
  1. 1.
    If $x_{i}<\bar{x}_{i}$, choose $y$ s.t. $y<x_{i}<t<\bar{x}_{i}$.
  2. 2.
    If $x_{i}>\bar{x}_{i}$, choose $y$ s.t. $\bar{x}_{i}<y<x_{i}<t$.
  3. 3.
    If $x_{i}=\bar{x}_{i}$, choose $y=x_{i}=\bar{x}_{i}$.
The game is not uniformly diagonally secure in the sens of Prokopovych2014 neither.
@@
Let’s construct such a sequence of deviations $(x_{i}^{k})$ for our game $\{(g_{i},[A,B])_{i=1}^{N}\}$. For any $x_{i}\in[A,B]$, the set of discontinuities $\mathbb{D}_{i}(x_{i})$ is the union of the following hypersurfaces:
@@
However, we get a problem at ties situation: If $\mathbf{x}_{-i}\in\mathbb{T}_{i}(x_{i})$, one can check $x_{i}^{k}\uparrow x_{i}$ is not a good choice when $x_{i}<E$. To bypass this obstacle, one could define the game $\{(\tilde{g}_{i},[A,B])_{i=1}^{N}\}$ where each payoff function $\tilde{g}_{i}$ has zero value on the ties. One can check for $x_{i}^{k}\downarrow x_{i}$ and $\mathbf{x}_{-i}\in\mathbb{T}_{i}(x_{i})$ that
@@
Theorem 5.4.
The game $\{(\tilde{g}_{i},[A,B])_{i=1}^{N}\}$ is diagonally payoff secure.
@@
Diagonal payoff security together with reciprocal upper semicontinuity implies diagonal better–reply security (Reny1999, Prop. 4.2). And one can try to deduce the existence of a symmetric Nash equilibrium in mixed strategies for this version of the game. However, the game $\{(\tilde{g}_{i},S_{i})_{i=1}^{N}\}$ is not reciprocal upper semicontinuous.
The auxiliary game $\{(\tilde{g}_{i},S_{i})_{i=1}^{N}\}$ resolves ties by assigning a null payoff at tie profiles. This corresponds to an exogenously specified tie-breaking rule, introduced solely for analytical convenience and without affecting payoffs outside a null set of strategy profiles.
A more general approach to payoff discontinuities induced by ties consists in allowing endogenous tie-breaking rules, whereby the resolution of ties is not fixed ex ante but is instead selected as part of the equilibrium outcome. Existence results for such games are provided by SimonZame1990, who study games with endogenous sharing rules and establish equilibrium existence under weak continuity requirements.
@@
Every compact metric game admits a mixed sharing rule solution.